# Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective

### v2

@@ 4 Energy-like metrics: what can go wrong
One important limitation of these metrics is that they indicate oversmoothing only in the limit of infinitely many layers, when their values converge exactly to zero. Since they measure a form of absolute distance, a small but nonzero value does not provide any meaningful information. On the other hand, convergence to zero corresponds to either perfect alignment with the dominant eigenspace or the collapse of the feature representation matrix to the all-zero matrix. While the former is a symptom of oversmoothing, the latter does not necessarily imply oversmoothing. Moreover, this convergence property requires the weights to be bounded. However, in practical cases, performance degradation is observed even in relatively shallow networks, far from being infinitely deep, and with arbitrarily large (or small) weight magnitudes. This aligns with our intuition and what occurs in the linear case. Indeed, for a linear GCN, even when the features $X^{(l)}$ grow to infinity as $l\to\infty$, one observes that $X^{(l)}$ becomes dominated by the dominant eigenspace of $A$, even for finite and possibly small values of $l$, depending on the spectral gap of the graph.
More precisely, the following theorem holds:
@@
Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN. Let $\lambda_{1},\lambda_{2}$ be the largest and second-largest eigenvalues (in modulus) of $A$, respectively. Assume the weights $\{W^{(l)}\}_{l=1}^{\infty}$ are randomly sampled from i.i.d. random variables with distribution $\nu$ such that
[…]
@@
with a linear rate of convergence $O(|\lambda_{2}/\lambda_{1}|^{l})$.
In particular, the theorem above implies that for a large spectral gap $|\lambda_{2}/\lambda_{1}|\ll 1$, $X^{(l)}$ is predominantly of rank one, namely
$X^{(l)}=\lambda_{1}^{l}\Big{(}uv^{T}+R(l)\Big{)}$
for some $v$, with $R(l)\sim O(|\lambda_{2}/\lambda_{1}|)^{l}$ and thus converging to 00 as $l\to\infty$. This results in weakly expressive feature representations, independently of the magnitude of the feature weights. This phenomenon can be effectively captured by measuring the rank of $X^{(l)}$, whereas Dirichlet-like energy measures may fail to detect it.
@@ 5 The Rank as a Measure of Oversmoothing
Inspired by the behaviour observed in the linear case, we argue that measuring the rank of feature representations provides a more effective way to quantify oversmoothing, in alignment with recent work on oversmoothing (Guo et al., 2023). However, since the rank of a matrix is defined as the number of nonzero singular values, it is a discrete function and thus not suitable as a measure. A viable alternative is to use a continuous relaxation that closely approximates the rank itself.
@@
Examples of possible continuous approximations of the rank include the numerical rank, the stable rank, and the effective rank (Roy & Vetterli, 2007; Rudelson & Vershynin, 2006; Arora et al., 2019). The stable rank is defined as $\text{StableRank}(X)=\|X\|_{*}^{2}/\|X\|_{F}^{2}$, where $\|X\|_{*}=\sum_{i}\sigma_{i}$ is the nuclear norm. The numerical rank is given by $\text{NumRank}(X)=\|X\|_{F}^{2}/\|X\|_{2}^{2}$. Finally, given the singular values $\sigma_{1}>\sigma_{2}>\dots>\sigma_{\min\{N,d\}}$ of $X$, the effective rank is defined as
$\text{Erank}(X)=\exp\left(-\textstyle{\sum_{k}p_{k}\log p_{k}}\right),$ (10)
@@
where $p_{k}=\sigma_{k}/\sum_{i}\sigma_{i}$ is the $k$-th normalized singular value. These rank relaxation measures exhibit similar empirical behaviour as shown by our numerical evaluation in Section 6.
@@
In practice, measuring oversmoothing in terms of a continuous approximation of the rank is a reasonable approach that helps address the limitations of Dirichlet-like measures. Specifically, it offers the following advantages: (a) it is scale-invariant, meaning it remains informative even when the feature matrix converges to zero or explodes to infinity; (b) it does not rely on a fixed, predetermined eigenspace but instead captures convergence of the feature matrix toward an arbitrary lower-dimensional subspace; (c) it allows for the detection of oversmoothing in shallow networks without requiring exact convergence to rank one—since a small value of, say, the effective rank, directly implies that the feature representations are low-rank, which in turn suggests a potentially suboptimal network architecture.
@@
In Table 1, we present a toy example illustrating that classical oversmoothing metrics fail to correctly capture oversmoothing unless the features are perfectly aligned. This observation implies that these metrics can quantify oversmoothing only when the rank of the feature matrix converges exactly to one. In contrast, continuous rank functions provide a more reliable measure of approximate feature alignment. Later, in Figure 1, we demonstrate that the same phenomenon occurs in GNNs trained on real datasets, where exact feature alignment is rare. In such cases, classical metrics remain roughly constant, whereas the rank decreases, coinciding with a sharp drop in GNN accuracy.
Table 1: A set of toy scenarios depicting the behaviour of oversmoothing metrics. Each plot contains 50 nodes, each with two features plotted as the x and y axis. # 1: all features are of the same value; # 2: all features are assumed to align with the same vector exactly; # 3: shifts one (red) point in # 2 in y-direction; # 4: samples all points from a uniform distribution. MAD (see Section 6) and $E_{\mathrm{Dir}}$ give false negative signals in # 3 while the features are considered oversmoothing by definition. $E_{\mathrm{Proj}}$ can hardly differentiate between # 3 and # 4, and is thus not robust in quantifying oversmoothing. To compute $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$ the first feature was considered in place of $u$ in (8) and (9).
@@
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 5.1 Theoretical Analysis of Rank Decay
In this section, we provide an analytical study proving the decrease of the numerical rank for a broad class of graph neural network architectures under the assumption that the weight matrices are entrywise nonnegative. While this is a somewhat restrictive setting, our result is the first theoretical proof that oversmoothing occurs independently of the weight (and thus feature) magnitude.
We begin with several useful observations. Let $u$ be the dominant eigenvector of $A$ corresponding to $\lambda_{1}$ and satisfying $\|u\|=1$. Consider the projection matrix $\mathcal{P}=uu^{\top}$. Given a matrix $X$, we can decompose it as $X=\mathcal{P}X+(I-\mathcal{P})X$. Since $u$ is a unit vector, it follows that $\|\mathcal{P}\|_{2}=1$, and therefore,
@@
Moreover, since $\mathcal{P}X$ and $(I-\mathcal{P})X$ are orthogonal with respect to the Frobenius inner product, we have $\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}=\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}$. Thus, we obtain the following bound for the numerical rank of $X$:
$\text{NumRank}(X)$ $=\frac{\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}$ (12)
$=\frac{\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}$
$\leq 1+\frac{\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}},$
where we used (11) and the fact that $\|\mathcal{P}X\|_{F}=\|\mathcal{P}X\|_{2}$ since $\mathcal{P}X$ is a rank-one matrix.
@@
Consider a linear GCN of the form $X^{(l+1)}=AX^{(l)}W^{(l)}$, where $A$ has a simple dominant eigenvalue $\lambda_{1}$ satisfying $|\lambda_{1}|\geq|\lambda_{2}|$.
We have already noted that $\|X\|_{2}\geq\|\mathcal{P}X\|_{2}$, meaning that the numerical rank converges to one if $\|(I-\mathcal{P})X\|_{F}/\|X\|_{2}$ decays to zero. This occurs whenever the features grow faster in the direction of the dominant eigenvector than in any other direction. As established in Theorem 4.1, this is almost surely the case in linear GNNs. As a direct consequence, we obtain the following result:
@@
Extending the result above to general GNNs with nonlinear activation functions is highly nontrivial. In the next section, we present our main theoretical result, which generalizes this analysis to a broader class of GNNs under the assumption of nonnegative weights.
5.2 The Nonnegative Setting
To study the case of networks with nonnegative weights, we make use of tools from the nonlinear Perron-Frobenius theory; we refer to (Lemmens & Nussbaum, 2012; Gautier et al., 2023) and the reference therein for further details.
We assume all the intermediate features to be in the positive open cone $\mathcal{K}:=\mathbb{R}_{+}^{N}=\{x\in\mathbb{R}^{N}\>|\>x_{i}>0\;\forall i=1,\dots,N\}$. On $\mathcal{K}$, it is possible to introduce the partial ordering
$x\leq_{\mathcal{K}}y\>(x\ll_{\mathcal{K}}y)\quad\text{iff}\quad y-x\in\bar{\mathcal{K}}\>(y-x\in\mathcal{K})$ (13)
where $\bar{\mathcal{K}}$ denotes the nonnegative closed cone. Given two points $x,y\in\mathcal{K}$ we write $M(x/y)=\max_{i}x_{i}/y_{i}$ and $m(x/y)=\min_{i}x_{i}/y_{i}$. Then, we can define the following Hilbert distance between any two points $x,y\in\mathcal{K}$,
$d_{H}(x,y)=\log(M(x/y)/m(x/y)).$
Note that $d_{H}$ is not a distance on $\mathcal{K}$, indeed $d_{H}(\alpha x,\beta y)=d_{H}(x,y)$ for any $x,y\in\mathcal{K}$ and $\alpha,\beta>0$. However it is a distance up to scaling; that is, it becomes a concrete distance whenever we restrict ourselves to a slice of the cone.
Now, it is well-known that any matrix $A$ such that $Ax\in\mathcal{K}$ for any $x\in\mathcal{K}$ is non-expansive with respect to the Hilbert distance, i.e.
@@ 5.2 The Nonnegative Setting
with $0\leq\beta\leq 1$. The last inequality is a consequence of the fact that $A$ preserves the ordering induced by the cone, i.e. if $x\geq_{\mathcal{K}}y$, then $Ax\geq_{\mathcal{K}}Ay$. In particular, we recall that (14) always hold for $\beta<1$ when $A$ is entry-wise strictly positive, as positive matrices map the whole $\bar{\mathcal{K}}$ in its interior, i.e. $x\geq_{\mathcal{K}}y$ implies $Ax>Ay$. In the next definition, we include all the nonnegative matrices that have the same behaviour as the positive matrices but only on a neighbourhood of a positive vector in $\mathcal{K}$.
Recall that any irreducible nonnegative matrix $A$ has a positive eigenvector $u$ corresponding to its dominant eigenvalue.
@@
A family of nonnegative irreducible matrices (possibly one single matrix) $\{A^{(l)}\}$ with the same dominant eigenvector $u>0$ is uniformly contractive with respect to $u$ if for any $C>0$ there exists some $\beta_{C}<1$ such that $d_{H}(A^{(l)}x,u)\leq\beta_{C}d_{H}(x,u)$ for all $l$ and $x\in\mathcal{K}$ such that $d_{H}(x,u)\leq C$.
The above definition is significantly weaker than asking for $A$ to be strictly positive. We show this with an illustrative example in Section A.3.
Consider now first the case of a linear function that can be represented in the form
$F(X)=AXW,$ (15)
where $X$ is a positive $N\times d$ matrix, $A$ and $W$ are nonnegative $N\times N$ and $d\times d$ matrices, respectively. Then, under mild assumptions, we can prove that the columns of $F(X)$ are closer to the dominant eigenvector $u$ of $A$ as compared to the columns of $X$.
@@
Let $F(X)=AXW$ with $A$ nonnegative and irreducible with dominant eigenvector $u\in\mathcal{K}$. Assume also $X$ to be striclty positve and $W$ nonnegative with $\min_{j}\max_{i}W_{ij}>0$. If $A$ is contractive with respect to $u$ and $C=\max_{i}d_{H}(X_{i},u)$ then
$\max_{i}d_{H}(F(X)_{i},u)\leq\beta_{C}\max_{i}d_{H}(X_{i},u),$
where $Y_{i}$ denotes the $i$-th column of $Y$, and $\beta_{C}\leq 1$.
@@
Subohomogeneity is a useful property and of practical utility. In fact, as discussed in, e.g. (Sittoni & Tudisco, 2024; Piotrowski et al., 2024), it is not difficult to verify that a broad range of activation functions commonly used in deep learning is subhomogeneous on $\mathcal{K}$. For this family of activation functions, Lemma 5.3 combined with arguments from nonlinear Perron–Frobenius theory yield the following main result
@@
Consider a positive GNN
$X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ (16)
with $X^{(0)}_{i}\in\mathcal{K}$ for any $i=1,\dots,d$ and $A^{(l)}$ and $W^{(l)}$ nonnegative for any $l$. Assume also that $u\in\mathcal{K}$ is the dominant eigenvector of all of the matrices $A^{(i)}$ and that it is also an eigenvector of $\sigma$. Then, if the matrices $\{A^{(i)}\}$ are uniformly contractive with respect to $u$ and $\min_{j}\max_{i}{W^{(l)}}_{ij}>0$ for any $l$, it holds
@@
We conclude with a formal investigation of the eigenpairs of activation functions that are entrywise subhomogeneous. Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R},\mathbb{R})$ that is subhomogeneous on $\mathbb{R}_{+}$. Then one can easily show that $\sigma$ is itself subhomogeneous on +$\mathbb{R}^{N}_{+}$. We have,
@@
As a consequence of the above result, we find that the numerical rank of the features collapses both for GCNs with LeakyReLU activation function (that is homogeneous) as well as for GATs with any kind of subhomogeneous activation function. In fact, the constant vector is always an eigenvector of an entriwise nonlinear map.
Table 2: The table lists the correlation coefficients of logarithm of metric values and the classification accuracies for GNNs of depth 2 - 24 layers and separately trained on Cora dataset. For Erank and NumRank, we subtract 1 so that both metrics converge to zero. The classification accuracy ratio are computed on 2-layered and 24-layered GNNs.
@@
LReLU + GCN -0.7871 0.6644 -0.8106 -0.8309 -0.2460 0.9724 0.5885 0.2689
Tanh + GCN 0.5243 0.9403 0.8610 0.9768 0.9734 0.9923 0.9784 0.1937
LReLU + GCN + Bias -0.9300 0.7505 -0.9414 -0.3642 -0.2828 0.9847 0.7833 0.2108
Tanh + GCN + Bias 0.4086 0.8950 0.7479 0.9026 0.8996 0.9926 0.9814 0.2133
LReLU + GCN + LayerNorm -0.9132 0.8445 -0.9424 0.5591 0.9753 0.9736 0.9736 0.4882
Tanh + GCN + LayerNorm -0.7119 0.2560 -0.2069 0.9716 0.8695 0.9576 0.9684 0.1885
LReLU + GCN + PairNorm -0.8165 0.3731 -0.8106 0.3885 0.4850 0.5597 0.6244 0.8252
Tanh + GCN + PairNorm -0.5963 0.0346 -0.5401 -0.3358 -0.0631 0.9838 0.9298 0.2902
LReLU + GAT -0.9189 0.6703 -0.9469 -0.6054 0.8251 0.9722 0.7612 0.2487
Tanh + GAT 0.8300 0.8501 0.8676 0.9066 0.8603 0.9600 0.9515 0.1767
Average correlation -0.3911 0.6279 -0.2722 0.2568 0.5296 0.9349 0.8541

### v3

@@ 4 Energy-like Metrics: What Can Go Wrong
One important limitation of these metrics is that they indicate oversmoothing only in the limit of infinitely many layers, when their values converge exactly to zero. Since they measure a form of absolute distance, a small but nonzero value does not provide any meaningful information. On the other hand, convergence to zero corresponds to either perfect alignment with the dominant eigenspace or the collapse of the feature representation matrix to the all-zero matrix. While the former is a symptom of oversmoothing, the latter does not necessarily imply oversmoothing. Moreover, this convergence property requires the weights to be bounded. However, in practical cases, performance degradation is observed even in relatively shallow networks, far from being infinitely deep, and with arbitrarily large (or small) weight magnitudes. This aligns with our intuition and what occurs in the linear case. Indeed, for a linear GCN, even when the features $X^{(l)}$ grow to infinity as $l\to\infty$, one observes that $X^{(l)}$ is dominated by the dominant eigenspace of $A$, even for finite and possibly small values of $l$, depending on the spectral gap of the graph. More precisely, the following theorem holds:
@@
with a linear rate of convergence $|\lambda_{2}/\lambda_{1}|$.
In particular, the theorem above implies that for a large spectral gap $|\lambda_{2}/\lambda_{1}|\ll 1$, $X^{(l)}$ is predominantly of rank one, namely $X^{(l)}=\lambda_{1}^{l}\Big{(}uv^{\top}+R(l)\Big{)}$ for some $v$, with $R(l)\sim O(|\lambda_{2}/\lambda_{1}|)^{l}$ and thus converging to 00 as $l\to\infty$. This results in weakly expressive feature representations, independently of the magnitude of the feature weights. This phenomenon can be effectively captured by measuring the rank of $X^{(l)}$, whereas Dirichlet-like energy measures may fail to detect it, this is e.g. the case when $\lambda_{1}>1$ having $X^{(l)}\rightarrow\infty$.
@@ 5 The Rank as a Measure of Oversmoothing
Inspired by the behaviour observed in the linear case, we argue that measuring the rank of feature representations provides a more effective way to quantify oversmoothing, in alignment with recent work on oversmoothing [14]. However, since the rank of a matrix is defined as the number of nonzero singular values, it is a discrete function and thus not suitable as a measure. A viable alternative is to use a continuous relaxation that closely approximates the rank itself.
@@
Examples of possible continuous approximations of the rank include the numerical rank, the stable rank, and the effective rank [31, 32, 1]. The stable rank is defined as $\text{StableRank}(X)=\|X\|_{*}^{2}/\|X\|_{F}^{2}$, where $\|X\|_{*}=\sum_{i}\sigma_{i}$ is the nuclear norm. The numerical rank is given by $\text{NumRank}(X)=\|X\|_{F}^{2}/\|X\|_{2}^{2}$. Finally, given the singular values $\sigma_{1}>\sigma_{2}>\dots>\sigma_{\min\{N,d\}}$ of $X$, the effective rank is defined as
$\text{Erank}(X)=\exp\left(-\textstyle{\sum_{k}p_{k}\log p_{k}}\right),$
@@
where $p_{k}=\sigma_{k}/\sum_{i}\sigma_{i}$ is the $k$-th normalized singular value. These rank relaxation measures exhibit similar empirical behaviour as shown in Section 6.
@@
In practice, measuring oversmoothing in terms of a continuous approximation of the rank is a reasonable approach that helps addressing the limitations of Dirichlet-like measures. Specifically, it offers the following advantages: (a) it is scale-invariant, meaning it remains informative even when the feature matrix converges to zero or explodes to infinity; (b) it does not rely on a fixed, predetermined eigenspace but instead captures convergence of the feature matrix toward an arbitrary lower-dimensional subspace; (c) it allows for the detection of oversmoothing in shallow networks without requiring exact convergence to rank one. A small value of the effective rank directly implies that the feature representations are low-rank, suggesting a potentially suboptimal network architecture.
@@
In Figure 1, we present a toy example illustrating that classical oversmoothing metrics fail to correctly capture oversmoothing unless the features are perfectly aligned. This observation implies that these metrics can quantify oversmoothing only when the rank of the feature matrix converges exactly to one. In contrast, continuous rank functions provide a more reliable measure of approximate feature alignment. Later, in Figure 2, we demonstrate that the same phenomenon occurs in GNNs trained on real datasets, where exact feature alignment is rare. In such cases, classical metrics remain roughly constant, whereas the rank decreases, coinciding with a sharp drop in GNN accuracy.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 1: Toy scenarios depicting the behaviour of oversmoothing metrics. Each plot contains 50 nodes, each with two features plotted on the x-y axis. The features are: # 1 of the same value; # 2 perfectly aligned with the same vector; # 3 aligned to the same vector except for one (red) point; # 4 sampled from a uniform distribution. MAD (Sec. 6) and $E_{\mathrm{Dir}}$ give false negative signals in # 3 although features are oversmoothing by definition. $E_{\mathrm{Proj}}$ can hardly differentiate between # 3 and # 4, and is thus not robust in quantifying oversmoothing. To compute $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$, the first feature was considered in place of $u$ in (1) and (2).
@@ 5.1 Theoretical Analysis of Rank Decay
In this section, we provide an analytical study proving the decrease of the numerical rank for a broad class of graph neural network architectures under the assumption that the weight matrices are entry-wise nonnegative. While this is a somewhat restrictive setting, our result is the first theoretical proof that oversmoothing occurs independently of the weight (and thus feature) magnitude.
We begin with several useful observations. Let $u$ be the dominant eigenvector of $A$ corresponding to $\lambda_{1}$ and satisfying $\|u\|=1$. Consider the projection $\mathcal{P}=uu^{\top}$. Given a matrix $X$, we can decompose it as $X=\mathcal{P}X+(I-\mathcal{P})X$. Since $u$ is a unit vector, it follows that $\|\mathcal{P}\|_{2}=1$, and therefore,
@@
Moreover, since $\mathcal{P}X$ and $(I-\mathcal{P})X$ are orthogonal with respect to the Frobenius inner product, we have $\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}=\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}$. Thus, we obtain the following bound:
$\text{NumRank}(X)=\frac{\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}=\frac{\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}},$ (4)
@@
Consider a linear GCN of the form $X^{(l+1)}=AX^{(l)}W^{(l)}$, where $A$ has a simple dominant eigenvalue $\lambda_{1}$ satisfying $|\lambda_{1}|\geq|\lambda_{2}|$. We have already noted that $\|X\|_{2}\geq\|\mathcal{P}X\|_{2}$, meaning that the numerical rank converges to one if $\|(I-\mathcal{P})X\|_{F}/\|X\|_{2}$ decays to zero. This occurs whenever the features grow faster in the direction of the dominant eigenvector than in any other direction. As established in Theorem 4.1, this is almost surely the case in linear GNNs. As a direct consequence, we obtain the following result:
@@
Extending the result above to general GNNs with nonlinear activation functions is highly nontrivial. In the next subsection, we present our main theoretical result, which generalizes this analysis to a broader class of GNNs under the assumption of nonnegative weights.
5.1.1 The Nonlinear Case
To study the case of networks with nonlinear activations, we make use of tools from the nonlinear Perron-Frobenius theory; we refer to [18, 12] and the reference therein for further details.
Assume all the intermediate features of the network to be in the positive open cone $\mathcal{K}:=\mathbb{R}_{+}^{N}=\{x\in\mathbb{R}^{N}\>|\>x_{i}>0\;\forall i=1,\dots,N\}$. On $\mathcal{K}$, it is possible to introduce the partial ordering
$x\leq_{\mathcal{K}}y\>(x\ll_{\mathcal{K}}y)\quad\text{iff}\quad y-x\in\bar{\mathcal{K}}\>(y-x\in\mathcal{K})$
where $\bar{\mathcal{K}}$ denotes the closed cone of nonnegative vectors. Given two points $x,y\in\mathcal{K}$ we write $M(x/y)=\max_{i}x_{i}/y_{i}$ and $m(x/y)=\min_{i}x_{i}/y_{i}$. Then, we can define the following Hilbert distance between any two points $x,y\in\mathcal{K}$, $d_{H}(x,y)=\log(M(x/y)/m(x/y)).$ Note that $d_{H}$ is not a distance on $\mathcal{K}$, indeed $d_{H}(\alpha x,\beta y)=d_{H}(x,y)$ for any $x,y\in\mathcal{K}$ and $\alpha,\beta>0$. However it is a distance up to scaling; that is, it becomes a concrete distance whenever we restrict ourselves to a slice of the cone. It is well-known that any matrix $A$ such that $Ax\in\mathcal{K}$ for any $x\in\mathcal{K}$ is non-expansive with respect to the Hilbert distance, i.e.
@@ 5.1.1 The Nonlinear Case
with $0\leq\beta\leq 1$. The last inequality is a consequence of the fact that $A$ preserves the ordering induced by the cone, i.e. if $x\geq_{\mathcal{K}}y$, then $Ax\geq_{\mathcal{K}}Ay$. In particular, we recall that (5) always hold for $\beta<1$ when $A$ is entry-wise strictly positive, as positive matrices map the whole $\bar{\mathcal{K}}$ in its interior, i.e. $x\geq_{\mathcal{K}}y$ implies $Ax>Ay$. In the next definition, we include all the nonnegative matrices that have the same behaviour as the positive matrices but only on a neighbourhood of a positive vector in $\mathcal{K}$. To this end, we recall that any irreducible nonnegative matrix $A$ has a positive eigenvector $u$ corresponding to its dominant eigenvalue.
@@
A family of nonnegative irreducible matrices $\{A^{(l)}\}$ with the same dominant eigenvector $u>0$ is uniformly contractive with respect to $u$ if for any $C>0$ there exists some $\beta_{C}<1$ such that $d_{H}(A^{(l)}x,u)\leq\beta_{C}d_{H}(x,u)$ for all $l$ and $x\in\mathcal{K}$ such that $d_{H}(x,u)\leq C$
Note that the above definition is significantly weaker than asking for $A$ to be strictly positive. We show this with an illustrative example in Section A.3.
Consider now first the case of a linear function that can be represented in the form $F(X)=AXW$, where $X$ is a positive $N\times d$ matrix, $A$ and $W$ are nonnegative $N\times N$ and $d\times d$ matrices, respectively. Then, under mild assumptions, we can prove that the columns of $F(X)$ are closer to the dominant eigenvector $u$ of $A$ as compared to the columns of $X$.
@@
Let $F(X)=AXW$ with $A$ nonnegative and irreducible with dominant eigenvector $u\in\mathcal{K}$. Assume also $X$ to be strictly positive and $W$ nonnegative with $\min_{j}\max_{i}W_{ij}>0$. If $A$ is contractive with respect to $u$ and $C=\max_{i}d_{H}(X_{i},u)$ then
$\max_{i}d_{H}(F(X)_{:,i},u)\leq\beta_{C}\max_{i}d_{H}(X_{:,i},u),$
where $Y_{:,i}$ denotes the $i$-th column of $Y$, and $\beta_{C}<1$.
@@
Subhomogeneity is a useful property and of practical utility. In fact, as discussed in  [38, 27], a broad range of activation functions commonly used in deep learning is subhomogeneous on $\mathcal{K}$. For this family of activation functions, Lemma 5.3 combined with arguments from nonlinear Perron–Frobenius theory yield the following main result
@@
Consider a GNN of the form $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ with $X^{(l)}_{:,i}\in\mathcal{K}$ for any $i=1,\dots,d$ and $A^{(l)}$ and $W^{(l)}$ nonnegative, for any $l$. Assume also that $u\in\mathcal{K}$ is the dominant eigenvector of all of the matrices $A^{(l)}$ and that it is also an eigenvector of $\sigma$. Then, if the matrices $\{A^{(l)}\}$ are uniformly contractive with respect to $u$ and $\min_{j}\max_{i}W^{(l)}_{ij}>0$ for any $l$, it holds
@@
We conclude with a formal investigation of the eigenpairs of activation functions that are entry-wise subhomogeneous. Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R},\mathbb{R})$ that is subhomogeneous on $\mathbb{R}_{+}$. Then one can easily show that $\sigma$ is itself subhomogeneous on +$\mathbb{R}^{N}_{+}$. We have,
@@
As a consequence of the above result, we find that the numerical rank of the features collapses to one both for GCNs with LeakyReLU activation function (the setting of [4]) as well as for GATs with any kind of subhomogeneous activation function (the setting of [45]), as the constant vector is always an eigenvector of an entry-wise nonlinear map. For GCNs with any other subhomogeneous activation functions, the convergence of the numerical rank to 1 as stated in Theorem 5.1 may not hold as discussed in Appendix D.
Dataset Model $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
@@
Cora GCN -0.7871 0.6644 -0.8106 -0.8309 -0.2460 0.9724 0.5885 0.2693
GAT -0.9189 0.6703 -0.9469 -0.6054 0.8251 0.9722 0.7612 0.2493
Citeseer GCN -0.8442 0.4350 -0.8913 -0.8667 -0.7169 0.9700 0.6795 0.4380
GAT -0.9576 0.0664 -0.9585 -0.9080 0.3722 0.9915 0.8047 0.4672
Pubmed GCN -0.9068 0.7006 -0.8508 -0.1109 0.6205 0.9464 0.9268 0.5225
GAT -0.8735 -0.3684 -0.8541 -0.4102 -0.3932 0.9270 0.9721 0.5564
Squirrel GCN -0.7774 0.4171 -0.7602 -0.3258 -0.8247 0.6316 0.9582 0.8457
GAT -0.6864 -0.5503 -0.7364 -0.7253 0.5002 0.8538 0.6840 0.7533
Chameleon GCN -0.9223 0.1504 -0.9163 -0.8201 -0.8809 0.9387 0.9014 0.6195
GAT -0.8721 0.1942 -0.9089 -0.8234 0.2803 0.9446 0.8799 0.6332
Amazon Ratings GCN -0.9297 0.8809 -0.9079 -0.3423 0.9201 0.9301 0.8049 0.8562
GAT -0.9388 0.5277 -0.9089 -0.1617 0.6545 0.9248 0.8764 0.8384
OGB-Arxiv GCN 0.7738 0.9194 0.5740 -0.2738 0.2822 0.9682 0.9091 0.0939
GAT -0.4097 0.9439 -0.7230 0.8985 0.8492 0.7740 0.9781 0.2310
Average correlation -0.7179 0.4036 -0.7571 -0.4504 0.1601 0.9103 0.8374
Table 1: Correlation between the classification accuracy and the logarithm of metric values on GNNs with LeakyReLU and of depths from 2 to 24 layers, separately trained on different homophilic (Cora, Citeseer, Pubmed), heterophilic (Squirrel, Chameleon, Amazon Ratings) and large-scale (OGB-Arxiv) datasets. For Erank and NumRank, we subtract 1 so that both metrics approach zero. The rightmost column reports the ratio of classification accuracy between GNNs with 2 and 24 layers. Some heterophilic datasets may be more resilient to the increasing network depth, in-line with observations from the literature, e.g. [13]. Additional results on other datasets and activation functions are presented in Appendix E, and with additional network components in Appendix F.